# The Host Galaxies of Radio AGN: New Views from Combining LoTSS and MaNGA Observations

### v1

@@
The role of radio mode active galactic nuclei (AGN) feedback on galaxy evolution is still under debate. In this study, we utilize a combination of radio continuum observations and optical integral field spectroscopic (IFS) data to explore the impact of radio AGN on the evolution of their host galaxies at both global and sub-galactic scales. We construct a comprehensive radio-IFS sample comprising 5578 galaxies with redshift by cross-matching the LOFAR Two-Metre Sky Survey (LoTSS) with the Mapping Nearby Galaxies at APO (MaNGA) survey. We revisit the tight linear radio continuum - star formation relation and quantify its intrinsic scatter, then use the relation to classify 616 radio-excess AGN with excessive radio luminosities over that expected from their star formation rate. Massive radio AGN host galaxies are predominantly quiescent systems, but the quenching level shows no correlation with the jet luminosity. The mass assembly histories derived from the stellar population synthesis model fitting agree with the cosmological simulations incorporating radio-mode AGN feedback models. We observe that radio AGN hosts grow faster than a control sample of galaxies matched in stellar mass, and the quenching age () is at larger lookback times than the typical radio jet age (). By stacking the spectra in different radial bins and comparing results for radio AGN hosts and their controls, we find emission line excess features in the nuclear region of radio AGN hosts. This indicates that radio AGN are ionizing the surrounding interstellar medium in the vicinity of the nucleus. We also find that ongoing star formation in the outer regions of the galaxy is weaker if a radio jet is detected. Our findings support the scenario that the observed present-day radio AGN activity is not responsible for the past quenching of their hosts, but may help the host galaxies maintain quiescence through ionizing and heating the surrounding gas.
@@
Much effort over the past decade, both in observations and simulations, has been devoted to the co-evolution of galaxies and their central supermassive black holes (SMBHs). One of the first pieces of evidence leading to this investigation was the tight correlation between the masses of SMBHs and the stellar velocity dispersion of their host galaxies and bulges (e.g., Magorrian et al. 1998). Additional indirect evidence supporting the co-evolution scenario is the observation that the cosmic growth history of SMBHs and the galaxy population is quite similar (e.g., Silverman et al. 2008). They both peak at redshift and decline rapidly after 1 (see Madau & Dickinson 2014, for a review). These observational results indicate that there are physical processes linking the growth of the SMBHs and their host galaxies, even though their sizes differ by orders of magnitude (pc versus kpc). However, the SMBH-galaxy co-evolution picture is still not fully understood in detail.
@@
The term AGN feedback describes the impact of AGN on the evolution of their host galaxies. A negative feedback scenario is widely favoured, in which the energy from accreting material can heat or eject the gas in the host galaxies, and then prevent star formation globally (e.g., Fabian 2012, and references therein). For most numerical and semi-analytic simulations, negative AGN feedback is crucial in massive halos to prevent galaxies from growing too massive (e.g., Di Matteo et al. 2005; Croton et al. 2006). On the other hand, the observed AGN-starburst connection in Type II Seyfert galaxies and infrared AGN indicate that positive AGN feedback may also exist (e.g., Heckman et al. 2004). The outflows and shocks from AGN may also create dense gas regions and boost star formation (e.g., Salomé et al. 2015; Schutte & Reines 2022). Alternatively, the correlation between AGN activity and star formation may arise because star formation and SMBH accretion share a common gas supply (e.g., Dai et al. 2018; Li et al. 2024). Recent theoretical models suggest that the dominant feedback mode transitions from positive to negative as redshift decreases (e.g., Silk et al. 2024).
@@
Radio AGN themselves have diverse observational properties. They are historically divided into Fanaroff & Riley (FR, Fanaroff & Riley 1974) types by radio morphology, where in general FR I, FR II, and FR0 (Garofalo & Singh 2019) represent ‘edge-brightened’, ‘edge-darkened’, and compact jets, respectively. Recent works suggested that the excitation properties may be more fundamentally connected to the SMBH accretion rate than the morphology (e.g., Best & Heckman 2012). Radio AGN with strong optical emission lines (e.g., [O III]5008) are classified as high-excitation radio galaxies (HERGs) with ongoing star formation activity in their hosts. The other population without strong emission lines are called low-excitation radio galaxies (LERGs).
@@
The new generation of radio telescopes and instruments, such as the LOw-Frequency ARray (LOFAR, van Haarlem et al. 2013), the Australian Square Kilometre Array Pathfinder (ASKAP, Hotan et al. 2021), MeerKAT (Jonas & MeerKAT Team 2016), etc., are now able to discover much fainter radio AGN. The LOFAR Two Meter Sky Survey (LoTSS, Shimwell et al. 2017) is one of the deepest ongoing radio continuum large-sky surveys, which aims to observe the whole northern sky at frequencies 120-168 MHz. Compared to NVSS and FIRST at 1.4 GHz, LoTSS observes at lower frequencies, which are more sensitive to old, diffuse synchrotron emission and less affected by the dust continuum thermal emission (e.g., Penney et al. 2020). For a extragalactic source with a typical radio spectra index, it is also about one order deeper than NVSS and FIRST, and sensitive enough to detect faint radio sources which are mainly star-forming galaxies (SFGs) at low redshifts (Williams et al. 2019; Hardcastle et al. 2023).
@@
In this paper, we will update the sample using the final MaNGA data release (Abdurro’uf et al. 2022), which includes 6000 galaxies falling within the LoTSS-DR2 sky footprint. This roughly doubles the size of the radio AGN sample compared to Mulcahey et al. (2022). Firstly, we aim to utilize this currently largest IFU-radio sample to provide an improved derivation of the radio continuum - SFR relation and its intrinsic scatter. Because MaNGA provides full-coverage H maps with dust attenuation information from the Balmer decrement, we can overcome the limitation of single fiber observations that only offer nuclear spectroscopic information for low-redshift galaxies. This also allows the radio continuum - SFR relation can be better constrained. Based on this relation, the excess radio luminosity compared to the prediction from its star formation rate is a physical and reliable criterion to classify radio-excess AGN (hereafter referred to as RDAGN). This relation to also be used to decompose the star formation and AGN contributions to the total radio flux, i.e., to calculate the jet luminosities (or upper limits).
@@
Throughout this paper, we convert all radio luminosity and related relations to rest-frame 144MHz assuming a power spectra index =-0.7 (Condon et al. 2002): . We adopt a flat cosmology with , , and . All stellar masses and star-formation rates are based on the Chabrier initial mass function (Chabrier 2003). When comparing literature results based on different IMFs, we follow Madau & Dickinson (2014) and set the factor (for stellar masses and star formation rates) as Salpeter : Kroupa : Chabrier = 1 : 0.67 : 0.63.
@@ 2.1 MaNGA
MaNGA is part of the Sloan Digital Sky Survey IV (SDSS-IV, Bundy et al. 2015), and is the largest completed IFU survey. The final data products are released in SDSS DR17 (Abdurro’uf et al. 2022), which consist of 10 000 unique galaxies with good quality IFU coverage out to at least 1.5 times the -band effective radius (). MaNGA galaxies are selected from the NASA-Sloan Atlas catalog111http://www.nsatlas.org/ (NSA, Blanton et al. 2011; Wake et al. 2017), and thus have measurements of many physical parameters based on the UV-optical photometric data, such as absolute magnitude, effective radius, position angle, etc. We adopt the global stellar mass from the NSA catalog to avoid having to correct for the fact that the MaNGA IFU sizes vary. The stellar masses are based on the stellar population synthesis model fitted from FUV to band photometry (Blanton & Roweis 2007). The redshift distribution of the sample is 0-0.15 and peaks at , while the stellar mass distribution is designed to have a flat distribution between (Wake et al. 2017). The IFU cubes have -band FWHM of 2.5 arcsec and cover 3600 - 10300 Å with a spectral resolution of (Smee et al. 2013). We note that the MaNGA sample will under-sample at the faint end (-band absolute magnitude larger than -19, Steffen et al. 2023), but this does not affect the main results in this paper since our radio AGN sample are dominated by massive galaxies (see Section 4.1).
@@ 2.2 LoTSS
In this paper, we use the source positions, flux density and spatial size from a value-added catalog with optical host galaxy identification information (hereafter LoTSS-VAC, Hardcastle et al. 2023). The host galaxy identification is based on a likelihood-ratio cross-match method described in Williams et al. (2019), as well as visual classifications from the citizen science project, Radio Galaxy Zoo: LOFAR222http://lofargalaxyzoo.nl/. The total flux density and morphological parameters have been examined, and where necessary corrected, in the host identification steps. This correction is necessary for deep radio observations in the MaNGA redshift range, because many radio galaxies have large sizes or diffuse emissions, creating large radio-optical offsets which will affect the traditional cross-matching method based on source separations. Furthermore, the erroneous segregation of a single radio galaxy into multiple components will result in the radio flux density being significantly underestimated. For the LoTSS-MaNGA matched sample used in this paper, the LoTSS-VAC recovers 44 radio galaxies missed in the raw PyBDSF catalog and corrects the flux density of 157 galaxies, which means that the raw catalog would have lost 8% accuracy.
@@ 2.3 LoTSS-MaNGA parent sample
Our analysis sample includes all MaNGA galaxies (to ensure a large enough IFU field of view) with LoTSS observations. Galaxies outside the LoTSS sky coverage or in the bad-quality sky areas are excluded based on the mosaic images, leaving us with 5687 MaNGA galaxies covered by LoTSS DR2. We match the MaNGA galaxy positions with the galaxy positions in LoTSS-VAC (i.e., the positions for the identified host galaxy, not the radio positions) within a radius of 9 arcsec (1.5beamsize and smaller than MaNGA IFU). This process results in 2655 detections (47%) with radio flux density and size measurements. We also visually check all the radio and optical images to exclude 109 (2%) galaxies with severe background source blending or having insufficient MaNGA coverage (IFU coverage is too small). For 2923 (51%) non-detections, we measure the median RMS flux within a 12 arcsec diameter aperture around the MaNGA positions in LoTSS images. We adopt 8 times the RMS flux density in one beam () as a secure flux density upper limit for each non-detection, since the catalog source completeness at this cut level reaches more than 95% (Hale et al. 2024).
@@
We calculate the rest-frame 144MHz luminosity or upper limit using the formula , where is the luminosity distance, is the galaxy redshift from MaNGA observation, and is assumed radio power-law index (-0.7, see Section 1). is the total flux density from LoTSS-VAC for detections, or 8 for non-detections. This rest-frame conversion only correct 2% of the value at MaNGA’s peak redshift , which is smaller than the level of most flux density errors, thus does not affect our analysis and results.
@@
Fig. 1 shows how the radio luminosity detection threshold varies with redshift for our sample of MaNGA galaxies with LoTSS-DR2 overlap. Blue dots represent LoTSS-detected galaxies and red upper limits are non-detections. The rest-frame 144MHz luminosity varies between . LoTSS DR2 reaches 90% completeness at 0.8 mJy (Shimwell et al. 2022), resulting a limit of at . We note that the exact limiting sensitivity also depends on the sky position, which causes the dispersion for the upper limits in Fig. 1. As a comparison, 72% and 41% of LoTSS detections will fall below the detection limits of NVSS (dashed line, Condon et al. 1998) and FIRST (dash-dotted line, Becker et al. 1995), respectively.
@@ 3.1 The -SFR relation
As mentioned in Section 1, synchrotron continuum luminosity is tightly correlated with the star formation rate in SFGs. Using LoTSS data, the radio-SFR relation has been investigated based on different SFR tracers and over different spatial scales. For example, Gürkan et al. (2018), Smith et al. (2021), and Best et al. (2023) used SFRs from SED-fitting for SFG populations out to , while Heesen et al. (2022) used SFRs converted from total infrared luminosity for resolved local group SFGs. Shenoy et al. (in prep) used both SED-fitted SFRs as well as integrated H derived SFRs. Up to now, there has been no relation derived from IFU-based H SFRs, despite the higher accuracy of this method. Taking advantage of the high-quality H maps from MaNGA, we derive and examine this radio-SFR relation for our LoTSS-MaNGA sample, and then use this relation to select radio-excess AGN (or RDAGN for short).
The details of how we derive this relation can be found in Appendix A. In brief, we select a pure star-forming galaxy sample without AGN contamination of both the H and radio continuum luminosities in two steps. Firstly we use the widely known emission line diagnostic diagrams (e.g., Kauffmann et al. 2003) to select galaxies where the H emission is dominated by H II regions. The emission lines used in this classification are measured from the MaNGA spectra. We also perform visual classification of the LoTSS images of these star-forming galaxies and remove those with jet-like or extended radio emission that is clearly larger than the galaxy. Our -SFR relation is then derived by performing a linear fit to the rest-frame 144MHz luminosities and SFRs, inspired by the efforts in Gürkan et al. (2018). The intrinsic errors of this relation are determined using the emcee (Foreman-Mackey et al. 2013) Markov chain Monte Carlo (MCMC) method.
@@
Fig. 2 shows that our results generally agree with previous low redshift relations based on different samples. For comparison, Brown et al. (2017) (blue line) used shallow TGSS sample with SFRs from single fiber H; Smith et al. (2021) (cyan line) and Best et al. (2023) (orange line) used a sample with deep radio observations and SFRs from multi-wavelength SED fitting; and Heesen et al. (2022) (green line) used a sample of local galaxies with SFRs measured from total infrared luminosities. Our results suggest an intrinsic scatter of dex in this relation, consistent with other LoTSS samples (e.g., Smith et al. 2021; Cochrane et al. 2023). The slope is larger than one, which supports the prediction of calorimetry models (e.g., Chi & Wolfendale 1990). This relation was found to also depend on the stellar mass (e.g., Smith et al. 2021; Heesen et al. 2022, Shenoy et al. in prep). Since the mass dependence is weak and more significant in low-mass systems (Das et al. 2024), we do not include the mass parameter in our fitting.

### v3

@@
The role of radio mode active galactic nuclei (AGN) feedback on galaxy evolution is still under debate. In this study, we utilize a combination of radio continuum observations and optical integral field spectroscopic (IFS) data to explore the impact of radio AGN on the evolution of their host galaxies at both global and sub-galactic scales. We construct a comprehensive radio-IFS sample comprising 5548 galaxies with redshift by cross-matching the LOFAR Two-Metre Sky Survey (LoTSS) with the Mapping Nearby Galaxies at APO (MaNGA) survey. We revisit the tight linear radio continuum - star formation relation and quantify its intrinsic scatter, then use the relation to classify 616 radio-excess AGN with excessive radio luminosities over that expected from their star formation rate. Massive radio AGN host galaxies are predominantly quiescent systems, but the quenching level shows no correlation with the jet luminosity. The mass assembly histories derived from the stellar population synthesis model fitting agree with the cosmological simulations incorporating radio-mode AGN feedback models. We observe that radio AGN hosts grow faster than a control sample of galaxies matched in stellar mass, and the quenching age () is at larger lookback times than the typical radio jet age (). By stacking the spectra in different radial bins and comparing results for radio AGN hosts and their controls, we find emission line excess features in the nuclear region of radio AGN hosts. This excess is more prominent in low-luminosity, low-mass, and compact radio AGN. The [N II]/ ratios of the excessive emission line indicate that radio AGN or related jets are ionizing the surrounding interstellar medium in the vicinity of the nucleus. Our results support the scenario that the observed present-day radio AGN activity may help their host galaxies maintain quiescence through gas ionization and heating, but it is not responsible for the past quenching of their hosts.
@@
Much effort over the past decade, both in observations and simulations, has been devoted to the co-evolution of galaxies and their central supermassive black holes (SMBHs). One of the first pieces of evidence leading to this investigation was the tight correlation between the masses of SMBHs and the stellar velocity dispersion of their host galaxy bulges (e.g., Magorrian et al. 1998; Ferrarese & Merritt 2000). Additional indirect evidence supporting the co-evolution scenario is the observation that the cosmic growth history of SMBHs and the galaxy population is quite similar (e.g., Silverman et al. 2008). They both peak at redshift and decline rapidly after 1 (see Madau & Dickinson 2014, for a review). These observational results indicate that there are physical processes linking the growth of the SMBHs and their host galaxies, even though their sizes differ by orders of magnitude (pc versus kpc). However, the SMBH-galaxy co-evolution picture is still not fully understood in detail.
@@
The term AGN feedback describes the impact of AGN on the evolution of their host galaxies. A negative feedback scenario is widely favoured, in which the energy from accreting material can heat or eject the gas in the host galaxies, and then prevent star formation globally (e.g., Fabian 2012, and references therein). For most numerical and semi-analytic simulations, negative AGN feedback is crucial in massive halos to prevent galaxies from growing too massive (e.g., Di Matteo et al. 2005; Croton et al. 2006). On the other hand, the observed AGN-starburst connection in Type II Seyfert galaxies and infrared AGN indicate that positive AGN feedback may also exist (e.g., Heckman et al. 2004). The outflows and shocks from AGN may also create dense gas regions and boost star formation (e.g., Salomé et al. 2015; Schutte & Reines 2022). Alternatively, the correlation between AGN activity and star formation may arise because star formation and SMBH accretion share a common gas supply (e.g., Dai et al. 2018; Li et al. 2024). Recent theoretical models suggest that the dominant feedback mode transitions from positive to negative as redshift decreases (e.g., Magliocchetti 2022; Silk et al. 2024).
@@
Radio AGN themselves have diverse observational properties. They are historically divided into Fanaroff & Riley (FR, Fanaroff & Riley 1974) types by radio morphology, where in general FR I, FR II, and FR0 (Garofalo & Singh 2019) represent ‘edge-darkened’, ‘edge-brightened’, and compact jets, respectively. Recent works suggested that the excitation properties may be more fundamentally connected to the SMBH accretion rate than the morphology (e.g., Best & Heckman 2012). Radio AGN with strong optical emission lines (e.g., [O III]5008) are classified as high-excitation radio galaxies (HERGs) with ongoing star formation activity in their hosts. The other population without strong emission lines are called low-excitation radio galaxies (LERGs).
@@
The new generation of radio telescopes and instruments, such as the LOw-Frequency ARray (LOFAR, van Haarlem et al. 2013), the Australian Square Kilometre Array Pathfinder (ASKAP, Hotan et al. 2021), MeerKAT (Jonas & MeerKAT Team 2016), etc., are now able to discover much fainter radio AGN. The LOFAR Two Meter Sky Survey (LoTSS, Shimwell et al. 2017) is one of the deepest ongoing radio continuum large-sky surveys, which aims to observe the whole northern sky at frequencies 120-168 MHz. Compared to NVSS and FIRST at 1.4 GHz, LoTSS observes at lower frequencies, which are more sensitive to old, diffuse synchrotron emission and less affected by the dust continuum thermal emission (e.g., Penney et al. 2020). For an extragalactic source with a typical radio spectra index, it is also about one order of magnitude deeper than NVSS and FIRST, and sensitive enough to detect faint radio sources which are mainly star-forming galaxies (SFGs) at low redshifts (Williams et al. 2019; Hardcastle et al. 2023).
@@
In this paper, we will update the sample using the final MaNGA data release (Abdurro’uf et al. 2022), which includes 6000 galaxies falling within the LoTSS-DR2 sky footprint. This roughly doubles the size of the radio AGN sample compared to Mulcahey et al. (2022). Firstly, we aim to utilize this currently largest IFU-radio sample to provide an improved derivation of the radio continuum - SFR relation and its intrinsic scatter. Because MaNGA provides full-coverage H maps with dust attenuation information from the Balmer decrement, we can overcome the limitation of single fiber observations that only offer nuclear spectroscopic information for low-redshift galaxies. This also allows the radio continuum - SFR relation to be better constrained. Based on this relation, the excess radio luminosity compared to the prediction from its star formation rate is a physical and reliable criterion to classify radio-excess AGN (hereafter referred to as RDAGN). This relation is also used to decompose the star formation and AGN contributions to the total radio flux, i.e., to calculate the jet luminosities (or upper limits).
@@
Throughout this paper, we convert all radio luminosity and related relations to rest-frame 144MHz assuming a radio spectral index =-0.7 (Condon et al. 2002): . We adopt a flat cosmology with , , and . All stellar masses and star-formation rates are based on the Chabrier initial mass function (Chabrier 2003). When comparing literature results based on different IMFs, we follow Madau & Dickinson (2014) and set the factor (for stellar masses and star formation rates) as Salpeter : Kroupa : Chabrier = 1 : 0.67 : 0.63.
@@ 2.1 MaNGA
MaNGA is part of the Sloan Digital Sky Survey IV (SDSS-IV, Bundy et al. 2015), and is the largest completed IFU survey. The final data products are released in SDSS DR17 (Abdurro’uf et al. 2022), which consist of 10 000 unique galaxies with good quality IFU coverage out to at least 1.5 times the -band effective radius (). MaNGA galaxies are selected from the NASA-Sloan Atlas catalog111http://www.nsatlas.org/ (NSA, Blanton et al. 2011; Wake et al. 2017), and thus have measurements of many physical parameters based on the UV-optical photometric data, such as absolute magnitude, effective radius, position angle, etc. We adopt the global stellar mass from the NSA catalog to avoid having to correct for the fact that the MaNGA IFU sizes vary. The stellar masses are based on the stellar population synthesis model fitted from FUV to band photometry (Blanton & Roweis 2007). The redshift distribution of the sample is 0-0.15 and peaks at , while the stellar mass distribution is designed to have a flat distribution between (Wake et al. 2017). The IFU cubes have -band FWHM of 2.5 arcsec and cover 3600 - 10300 Å with a spectral resolution of (Smee et al. 2013). We note that the MaNGA sample will under-sample at the faint end (-band absolute magnitude larger than -19, Steffen et al. 2023), but this does not affect the main results in this paper since our radio AGN sample is dominated by massive galaxies (see Section 4.1).
@@ 2.2 LoTSS
In this paper, we use the source positions, flux density and spatial size from a value-added catalog with optical host galaxy identification information (hereafter LoTSS-VAC, Hardcastle et al. 2023). The host galaxy identification is based on a likelihood-ratio cross-match method described in Williams et al. (2019), as well as visual classifications from the citizen science project, Radio Galaxy Zoo: LOFAR222http://lofargalaxyzoo.nl/. The optical and infrared information used for this cross-match is based on the DESI Legacy Imaging Surveys333https://legacysurvey.org/ (Dey et al. 2019) and the unWISE data (Schlafly et al. 2019). About 85% LoTSS detections have an optical or infrared identification (Hardcastle et al. 2023). The total flux density and morphological parameters have been examined, and where necessary corrected, in the host identification steps. This correction is necessary for deep radio observations in the MaNGA redshift range, because many radio galaxies have large sizes or diffuse emissions, creating large radio-optical offsets which will affect the traditional cross-matching method based on source separations. Furthermore, the erroneous association of a radio galaxy with only one of its multiple components will result in its the radio flux density being significantly underestimated. For the LoTSS-MaNGA matched sample used in this paper, the LoTSS-VAC recovers 44 radio galaxies missed in the raw PyBDSF catalog and corrects the flux density of 157 galaxies, which means that the raw catalog would have lost 8% accuracy.
@@ 2.3 LoTSS-MaNGA parent sample
Our analysis sample includes all MaNGA galaxies (to ensure a large enough IFU field of view) with LoTSS observations. Galaxies outside the LoTSS sky coverage or in the bad-quality sky areas are excluded based on the mosaic images, leaving us with 5687 MaNGA galaxies covered by LoTSS DR2. We match the MaNGA galaxy positions with the galaxy positions in LoTSS-VAC (i.e., the positions for the identified host galaxy, not the radio positions) within a radius of 9 arcsec (1.5beamsize and smaller than MaNGA IFU)444 This radius is safe enough since there is only one galaxy or one galaxy merger in a 20″ MaNGA cube, and can also include the sources with large SDSS-unWISE offsets (the FWHM of SDSS images and WISE W1 images are around 2.5″ and 6.6″, respectively). The actual matching offsets are usually far smaller than 9″, e.g. 90% detection have offsets less than 1″.. This process results in 2655 detections (47%) with radio flux density and size measurements. We also visually check all the radio and optical images to exclude 109 (2%) galaxies with severe background source blending or having insufficient MaNGA coverage (IFU coverage is too small). For 2923 (51%) non-detections, we measure the median RMS flux within a 9 arcsec radius aperture around the MaNGA positions in LoTSS images. We note that there are several non-detections close to luminous radio sources (e.g., MaNGA 8312-3704), and they have much higher RMS levels than sources in the normal fields due to artifacts around the strong sources. Thus, we consider the 2% of non-detections with the highest RMS values as outliers and exclude them from our final sample. We adopt 8 times the RMS flux density in one beam () as a secure flux density upper limit for each non-detection, since the catalog source completeness at this cut level reaches more than 95% (Hale et al. 2024).
@@
We calculate the rest-frame 144MHz luminosity or upper limit using the formula , where is the luminosity distance, is the galaxy redshift from MaNGA observation, and is assumed radio power-law index (-0.7, see Section 1). is the total flux density from LoTSS-VAC for detections, or 8 for non-detections. This rest-frame conversion only corrects 2% of the value at MaNGA’s peak redshift , which is smaller than the level of most flux density errors, thus does not affect our analysis and results.
@@
Fig. 1 shows how the radio luminosity detection threshold varies with redshift for our sample of MaNGA galaxies with LoTSS-DR2 overlap. Purple dots represent LoTSS-detected galaxies and grey upper limits are non-detections. The rest-frame 144MHz luminosity varies between . LoTSS DR2 reaches 90% completeness at 0.8 mJy (Shimwell et al. 2022), resulting in a limit of at . We note that the exact limiting sensitivity also depends on the sky position, which causes the dispersion for the upper limits in Fig. 1. A few non-detections significantly above the LoTSS limit are the sources affected by strong artifacts in the LoTSS images, which are usually caused by nearby luminous sources (Shimwell et al. 2022). We try to minimize the bias from these sources by taking into account both the upper limit and its error when selecting RDAGN and control samples. This is discussed in detail in Section 3.2 and 3.3.
The new LoTSS observations significantly improve the radio detection rate. As a comparison, 72% and 41% of LoTSS detections will fall below the detection limits of NVSS (dashed line in Fig. 1, Condon et al. 1998) and FIRST (dash-dotted line in Fig. 1, Becker et al. 1995), respectively. The LoTSS detection rate of MaNGA galaxies (Fig. 1 upper panel) stays approximately constant with redshift, but decreases noticeably at . This is due to the fact that the selection of MaNGA parent sample will constrain the apparent sizes of galaxies to not over-fill the IFUs, thus at lower redshifts the sample will be dominated by physically smaller, and hence lower mass, galaxies which tend to have faint radio emission.
@@ 3.1 The -SFR relation
As mentioned in Section 1, synchrotron continuum luminosity is tightly correlated with the star formation rate in SFGs. This relation can be used to select radio-excess AGN (or RDAGN for short). Using LoTSS data, the radio-SFR relation has been investigated based on different SFR tracers and over different spatial scales. For example, Gürkan et al. (2018), Smith et al. (2021), and Best et al. (2023) used SFRs from SED-fitting for SFG populations out to , while Heesen et al. (2022) used SFRs converted from total infrared luminosity for resolved local group SFGs. Shenoy et al. (in prep) used both SED-fitted SFRs as well as integrated H derived SFRs. Up to now, there has been no relation derived from IFU-based H SFRs, despite the higher accuracy of this method. For consistency, here we derive the -SFR relation from the LoTSS-MaNGA sample to help us select RDAGN later.
The details of how we derive this relation can be found in Appendix A. In brief, we select a pure star-forming galaxy sample without AGN contamination of both the H and radio continuum luminosities. Then the -SFR relation is derived by performing a linear fit to the rest-frame 144MHz luminosities and SFRs, inspired by the efforts in Gürkan et al. (2018). The intrinsic errors of this relation are determined using the emcee (Foreman-Mackey et al. 2013) Markov chain Monte Carlo (MCMC) method.
@@
Fig. 2 shows that our results generally agree with previous low redshift relations based on different samples. For comparison, Brown et al. (2017) (blue line) used shallow TGSS sample with SFRs from single fiber H; Smith et al. (2021) (cyan line) and Best et al. (2023) (orange line) used a sample with deep radio observations and SFRs from multi-wavelength SED fitting; and Heesen et al. (2022) (green line) used a sample of local galaxies with SFRs measured from total infrared luminosities. Our results suggest an intrinsic scatter of dex in this relation, consistent with other LoTSS samples (e.g., Smith et al. 2021; Cochrane et al. 2023). The slope is larger than one, which supports the prediction of calorimetry models (e.g., Chi & Wolfendale 1990).
This relation was found to also depend on the stellar mass (e.g., Smith et al. 2021; Heesen et al. 2022, Shenoy et al. in prep), but the mass dependence is weak and only significant in low-mass systems (Das et al. 2024) and at higher redshift (Delvecchio et al. 2021), thus we do not include the mass parameter in our fitting. Our fitting procedure also does not include the radio non-detections, which may result in a slightly more conservative cut for radio AGN classification. Though our result is reliable enough to classify bona fide radio AGN in massive galaxies (e.g. ), readers should be careful and test the mass dependence when applying it in lower mass samples.